OpenStudy (anonymous):

Solve the equation for y. 5x – 7y = 10

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

hey its me again :P

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Clear out any fractions by Multiplying every term by the bottom parts. Add or Subtract the same value from both sides. Divide every term by the same nonzero value. Combine Like Terms. Factoring. Expanding (the opposite of factoring) may also help.

OpenStudy (igreen):

Subtract 5x to both sides: -7y = -5x + 10 Now divide 7 to all terms: \(\dfrac{-7y}{-7} = \dfrac{-5x}{-7} + \dfrac{10}{-7}\) What do you get? @Chop-GTA5

OpenStudy (wmj259):

You start by subtracting 5x from both sides. and then dividing the whole thing by -7 to get just a single "y =" formula.

OpenStudy (anonymous):

Do you need help with that question

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

i will give you a blue medal

OpenStudy (anonymous):

so what the question

OpenStudy (anonymous):

Solve the equation for y. 5x – 7y = 10

OpenStudy (anonymous):

oh wait i got it

OpenStudy (igreen):

What is it?

OpenStudy (anonymous):

\[y=\frac{ 5 }{ 7 }x-\frac{ 10 }{ 7 }\]

OpenStudy (anonymous):

7y=-103x+2y=10

OpenStudy (igreen):

Yep, you got it! @Chop-GTA5

OpenStudy (anonymous):

okay thank you!

OpenStudy (igreen):

No problem.

OpenStudy (maryfitzgerald14):

thanks for helping me guys